# MultiNash-PF: A Particle Filtering Approach for Computing Multiple Local Generalized Nash Equilibria in Trajectory Games

### v2

@@ Abstract
Modern robotic systems frequently engage in complex multi-agent interactions, many of which are inherently multi-modal, meaning they can lead to multiple distinct outcomes. To interact effectively, robots must recognize the possible interaction modes and adapt to the one preferred by other agents. In this work, we propose an efficient algorithm for capturing the multimodality in multi-agent interactions. We leverage a game-theoretic planner to model interaction outcomes as equilibria where each equilibrium corresponds to a distinct interaction mode. We then develop an efficient algorithm to identify all the equilibria, allowing robots to reason about multiple interaction modes. More specifically, we formulate interactive planning as Constrained Potential Trajectory Games (CPTGs) and model interaction outcomes by local Generalized Nash equilibria (GNEs) of the game. CPTGs are a class of games for which a local GNE can be found by solving a single constrained optimal control problem where a potential function is minimized. We propose to integrate the potential game approach with implicit particle filtering, a sample-efficient method for non-convex trajectory optimization. We utilize implicit particle filtering to identify the coarse estimates of multiple local minimizers of the game’s potential function. MultiNash-PF then refines these estimates with optimization solvers, obtaining different local GNEs. We show through numerical simulations that MultiNash-PF reduces computation time by up to 50% compared to a baseline. We further demonstrate the effectiveness of our algorithm in real-world human-robot interaction scenarios, where it successfully accounts for the multi-modal nature of interactions and resolves potential conflicts in real-time.
@@ I Introduction
Modern-world robotics applications are inherently multi-agent in nature. For instance, autonomous cars on highways need to account for other vehicles while planning their motion. Most of the real-world interactions are multi-modal. For example, in the two-player interaction illustrated in Fig. 1, the agents need to move towards their goals while avoiding collisions. Here, two outcomes exist: both agents yielding to their right or left. If the two agents were to act independently, they could reach different outcomes, which would result in a collision. What trajectory each agent would prefer to pick depends on the convention they might be following. For example, in some cultures, in such settings, the convention is to yield to the right to avoid collisions while in other cultures, the convention is to yield to the left. Such conventions usually arise out of recurring coordination problems [1]. Therefore, autonomous agents need to be aware of the existence of multi-modality in the interactions for efficient coordination with people who follow different conventions. Some recent works have shown that computing such multiple trajectories is critical to aligning agents’ preferences over different outcomes [2] or helping coordinate agents’ actions via recommendation [3].
@@
Figure 1: Snapshot of an interaction between a robot and a human when they exchange positions. Using MultiNash-PF the robot can compute both modes of interaction. At $t=2$, the robot is uncertain about the preferred mode of the human since the human trajectory’s distance from both equilibria is almost the same. At $t=4$, the human trajectory is closer to the second GNE therefore, the robot chooses to follow the second GNE accordingly and avoid collision with the human.
@@
To formalize this problem, we will use dynamic game theory [4] which has shown to be powerful in developing interactive and socially complaints robots [5, 6, 7, 8]. We formulate a multi-agent interactive trajectory planning problem as a Constrained Potential Trajectory Game (CPTG). In CPTGs, the interaction outcomes are captured through generalized Nash equilibrium (GNE), which are constrained Nash equilibria. When multiple agents- such as robots and humans- interact, a GNE corresponds to a joint action such that no agent has incentives to deviate from its equilibrium actions unilaterally. Several recent works have shown great success in utilizing GNE for multi-agent motion planning in interactive domains [9, 10, 11, 12]. A key challenge in interactive domains, however, is the existence of multiple equilibria, many of which are equally valid. We argue that each equilibrium corresponds to a distinct mode of interaction. For instance, in Fig. 1, yielding to the right or left are both acceptable equilibria, representing different Nash equilibria. Consequently, we propose that in order to capture different interaction modes, robots must reason about multiple interaction equilibria and select the one that aligns with the convention adopted by other agents.
Computing different local GNEs requires an efficient way to solve games. In practice, computing even a single GNE is often challenging [13] because solving for GNE amounts to solving a set of coupled constrained optimal control problems. Typically, such a solution requires solving first-order KKT conditions of the underlying problems. Therefore, solving the GNE often reduces to addressing a set of complementarity conditions, a process known to be computationally demanding, especially for large-scale applications [14]. Furthermore, computing all the different local GNEs requires initializing the game solvers with many different initial guesses of the equilibrium trajectories from prior knowledge of the specific scenario. This further increases the required number of initial guesses for the game solvers to find all local GNEs. To our best knowledge, there is no principled approach to compute multiple local Nash equilibria in constrained trajectory games other than exhaustive random initializations of the game solvers.
Figure 2: The Nash equilibrium trajectories found by MultiNash-PF when two unicycle agents change their positions. First, we use the implicit particle filter to discover the coarse estimates of two different equilibria, and then utilize them as a warm starting for IPOPT to obtain the two Nash equilibrium trajectories.
@@
In this paper, we tackle this challenge by introducing MultiNash-PF, a novel framework designed to efficiently compute multiple local generalized Nash equilibria (GNEs) in constrained trajectory games. This, in turn, enables robots to recognize and reason about multiple interaction modes. Our approach integrates constrained potential games [15], a class of games that allow for the computation of local GNEs by solving a single constrained optimal control problem, with constraint-aware implicit particle filtering—a sample-efficient method recently applied to nonconvex trajectory optimization [16] to achieve efficient identification of multiple equilibria.
@@
Using CPTGs, we can compute a local GNE by solving a single constrained optimal control problem using off-the-shelf optimization solvers. Then, we use a constraint-aware implicit particle filtering method to identify multiple local minima of the potential minimization problem. This method eliminates the requirements of random initialization of the optimization solver by efficiently obtaining coarse estimates of different local GNEs. Next, we use the identified minima as initialization for an optimization solver to refine our solution. The solver then computes all the different local GNEs of the original game. Leveraging the sample efficiency of the implicit particle filtering method, MultiNash-PF can explore and identify the multiple local GNEs faster than the random initialization of optimization solvers. We further show the real-time application of MultiNashPF in a human-robot navigation setting where the robot can reason about the existence of different local GNEs and plan its motion autonomously based on the modality the human prefers.
Figure 3: MultiNash-PF identifies all the local GNE trajectories for a challenging two agent trajectory game where two agents swapped positions while avoiding an obstacle. In this scenario, the obstacle radius is greater than the collision avoidance radius of the two agents, leading to the discovery of 6 distinct Nash equilibria. MultiNash-PF effectively obtains coarse estimates of local GNEs through implicit particle filtering, which are then refined using IPOPT to obtain local GNEs.
@@ II Related Works
Game Theoretic Planning. One of the earliest works laying the foundation of game theoretic motion planning was [17]. Game theoretic motion planners that seek Stackelberg equilibria were proposed in [18, 7]. Iterative best response methods for computing Nash equilibria were developed in [19, 20]. Iterative linear quadratic approximations to compute Nash equilibria were introduced in [6]. ALGames was developed in [21] which is an augmented lagrangian-based solver for constrained dynamic games. A sequential linear-quadratic technique game-based technique to compute GNE was proposed in [22]. However, these methods are computationally expensive for real-time implementations as the number of agents increases. Potential games-based methods were developed later in [15, 23, 11, 12] to reduce the complexity of computing equilibria. However, the shortcomings of all these methods lie in the fact that they find only one Nash equilibrium of the underlying game and ignore the multi-modal nature of the problem that is crucial for coordination among agents. In [2], the authors recognize the existence of multiple equilibria and focus on inferring the equilibrium of other agents in the environment. However, the methodology relies on computing the different approximate local Nash by rollouts of randomly selected initial strategies, which we show in this work performs poorly compared to our method.
Multiple Local Solutions to Constrained Optimal Control. In general, optimal control problems are often formulated as nonconvex constrained optimization problems, which are solved through the conventional gradient-based approaches such as inter-point methods [24] and sequential quadratic programming [25]. However, given an initialization, they can only find a single local optimal solution. However, in the present context of identifying all potential local GNEs, the gradient-based method would require numerous initializations to find all the local solutions. Departing from gradient-based methods, a Bayesian inference approach has been proposed to solve model predictive control problem [26, 27, 28]. The approach formulates the optimal control problem from a Bayesian inference perspective and employs particle filtering as its estimation method. At its core, particle filtering utilizes sequential Monte Carlo sampling to perform state estimation for highly nonlinear systems [29, 30]. The efficiency and accuracy of the estimation largely depend on the quality of the sampled particles. Recently, in [31], an efficient particle filtering method was proposed based on the principle of implicit sampling.
@@ III Constrained Potential Trajectory Games
We formulate the multi-agent interactive trajectory planning problem as an $N$-agent constrained potential trajectory game (CPTG). Let $\mathcal{N}:=\{1,2,\ldots,N\}$ denote the set of agents’ indices. We assume that the game is played over a finite time horizon of $\tau\in\mathbb{N}$ time-steps, where $\mathbb{N}$ is the set of natural numbers.
Let $x^{i}_{t}\in\mathbb{R}^{n_{i}}$ and $u_{t}^{i}\in\mathbb{R}^{m_{i}}$ denote the state and control of the $i$-th agent at time $t$, respectively. We assume that the state of the $i$-th agent evolves as follows:
@@
In CPTGs, each agent has to satisfy some constraints that couple different agents’ states and controls. Let $g:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{c}$ with $c\in\mathbb{N}$ denote a continuously differentiable function that defines the constraints on the joint state $x_{t}$ and joint control $u_{t}$ at each time:
@@
where ${0}_{c}$ is a vector of zeros in $\mathbb{R}^{c}$. The aim of the constraint function $g$ is to capture individual agents’ strict preferences, such as each agent avoiding obstacles, or avoiding collision with other agents. It should be noted that coupling amongst agents’ decisions is captured through constraints in CPTGs.
@@
where $\hat{x}_{t}^{i}\in\mathbb{R}^{n_{i}}$ is the reference state for agent $i$ at time $t$. Typically, the reference states for each agent is a path of minimum cost they would follow in the absence of other agents in the environment, e.g., a straight line connecting the start and goal location. Furthermore, we denote the joint reference state of all agents at time $t$ as $\hat{x}_{t}$. We use $\mathop{\rm blkdiag}(A_{1},\ldots,A_{n})$ to denote the block-diagonal of the matrices $A_{1},\ldots,A_{n}$. Let $a_{0:\tau}:=\{a_{0},\ldots,a_{\tau}\}$ denote the collection of vectors $a_{t}$’s over the entire horizon $0\leq t\leq\tau$. We use $\{{x}_{0:\tau}^{i},{u}_{0:\tau}^{i}\}_{i=1}^{N}$ to denote joint trajectories of all $N$ agents across all times. Note that using the joint notation, we can equivalently write the joint trajectories of all agents as $\{{x}_{0:\tau},{u}_{0:\tau}\}$.
@@ III-A Local Generalized Nash Equilibrium Trajectories
Due to the interactive nature of the problem, generally, it is not possible for all agents to minimize their costs simultaneously while satisfying the constraints. Therefore, the interaction outcome is best captured by local GNE trajectories of the underlying trajectory game. With dynamics as (1) and constraints as (2), we denote the game by
@@
A set of joint trajectories $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}_{i=1}^{N}$ is a local GNE trajectory for $\mathcal{G}$ if for all agent $i\in\mathcal{N}$, agent i’s trajectory $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}$ is an optimal solution of the following optimization problem withing a local neighborhood $\mathcal{D}\left(\{{x}^{\star}_{0:\tau},{u}^{\star}_{0:\tau}\},\epsilon\right)$ for some $\epsilon>0$:
@@
Intuitively, at a local GNE, no agent can reduce their cost function by independently changing their trajectory to any alternative feasible trajectory within a local neighborhood of the joint equilibrium trajectory, $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}_{i=1}^{N}$. Note that in CPTGs, the agents’ trajectories are coupled through their joint constraints. It is important to highlight that solving (1) requires solving a set of $N$ coupled constrained optimal control problems, which is often computationally expensive. Therefore, finding all the local GNEs of the game can be time-consuming when using traditional game theoretic methods that aim to solve coupled optimal control problems.
However, an important property of CPTGs is that a GNE of the original game can be computed by optimizing a single constrained optimal control problem called Potential, as described in [11]. This reduces the computational cost of computing the GNEs of the game. In the following, we will prove that game $\mathcal{G}$ can be expressed as a constrained trajectory potential game. The following theorem shows that any local minimizer of such a potential function is a local generalized Nash equilibrium trajectory.
@@
A set of joint trajectories $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}_{i=1}^{N}$ is a local GNE trajectory if within a local neighborhood $\mathcal{D}\left(\{{x}^{\star}_{0:\tau},{u}^{\star}_{0:\tau}\},\epsilon\right)$, it is an optimal solution of
@@
Furthermore, using the joint notations, we can re-write the sums in (6) as (1). Therefore, we have proven that $\mathcal{G}$ is a CPTG with potential (1). Hence, using Theorem 2 from [11], any local solution of (1) will also be a local GNE of $\mathcal{G}$. ∎
Using Theorem 1, any local solutions of (1) will also be local GNE trajectories. We would like to acknowledge that finding local solutions of (1) will not necessarily give us all the GNEs of the original as Theorem 1 provides only a sufficient condition for computing local GNEs. Furthermore, it is known that the problem of finding all Nash equilibria is a PPAD-complete problem even in the case of static games [13]. However, for all practical purposes in multi-agent navigation, finding local solutions of (1) has proven to be enough for us.
The conventional approach to solving the optimization (1) relies on techniques such as interior point methods [24] or sequential quadratic programming [25]. Naively, the local solutions can be obtained by exhaustive initialization of random trajectories to an optimization solver, such as IPOPT [24], hoping to explore the solution space of local Nash equilibria. However, randomly exploring the nonconvex optimization landscape via gradient-based methods is computationally prohibitive and inefficient, as the same local equilibria can be arrived at from multiple random initializations. To address this, we utilize a Bayesian inference approach in the following section to efficiently recover all local minima of (1).

### v3

@@ Abstract
Modern robotic systems frequently engage in complex multi-agent interactions, many of which are inherently multi-modal, i.e., they can lead to multiple distinct outcomes. To interact effectively, robots must recognize the possible interaction modes and adapt to the one preferred by other agents. In this work, we propose MultiNash-PF, an efficient algorithm for capturing the multimodality in multi-agent interactions. We model interaction outcomes as equilibria of a game-theoretic planner, where each equilibrium corresponds to a distinct interaction mode. Our framework formulates interactive planning as Constrained Potential Trajectory Games (CPTGs), in which local Generalized Nash Equilibria (GNEs) represent plausible interaction outcomes. We propose to integrate the potential game approach with implicit particle filtering, a sample-efficient method for non-convex trajectory optimization. We utilize implicit particle filtering to identify the coarse estimates of multiple local minimizers of the game’s potential function. MultiNash-PF then refines these estimates with optimization solvers, obtaining different local GNEs. We show through numerical simulations that MultiNash-PF reduces computation time by up to 50% compared to a baseline. We further demonstrate the effectiveness of our algorithm in real-world human-robot interaction scenarios, where it successfully accounts for the multi-modal nature of interactions and resolves potential conflicts in real-time.
@@ I Introduction
Many real-world robotic interactions are inherently multi-modal. For example, in the two-player interaction illustrated in Fig. 1, the agents need to move towards their goals while avoiding collisions. Here, two outcomes exist: both agents yielding to their right or left. If the two agents act independently, they might choose conflicting actions. For example, one yields to the left and the other to the right, which can potentially result in a collision. What trajectory each agent would prefer to pick depends on the convention they might be following. For example, in some cultures, in such settings, the convention is to yield to the right to avoid collisions, while in other cultures, the convention is to yield to the left. Such conventions typically emerge from repeated coordination problems [1], and autonomous agents must recognize them and account for the multi-modality of interactions to coordinate with humans effectively. Some recent works have shown that reasoning about the multi-modality of interactions is critical to aligning agents’ preferences over different outcomes [2] or helping coordinate agents’ actions via recommendation [3]. In this paper, we address this challenge by proposing MultiNash-PF, an algorithm that efficiently identifies multiple likely interaction outcomes in multi-agent domains and enables agents to adapt their plans accordingly.
@@
Figure 1: Snapshot of an interaction between a robot and a human when they exchange positions. Using MultiNash-PF, the robot can compute both modes of interaction. At $t=2$, the robot is uncertain about the preferred mode of the human since the human trajectory’s distance from both equilibria is almost the same. At $t=4$, the human trajectory is closer to the yield-right GNE; therefore, the robot chooses the corresponding GNE and avoids collision with the human.
@@
We formulate the problem of computing multi-agent interaction outcomes as a multi-agent trajectory planning problem. To formalize interactions, we use dynamic game theory [4], which has been shown to be powerful in developing interactive and socially compliant robots [5, 6, 7, 8] where it has been shown that the interaction outcomes are captured through generalized Nash equilibrium (GNE). When multiple agents interact, a GNE corresponds to a joint action such that no agent has incentives to deviate from its equilibrium actions unilaterally. Several recent works have shown great success in utilizing GNE for modeling interaction outcomes in interactive domains [9, 10, 11, 12]. However, there typically exist multiple equilibria in interaction games, many of which are equally valid. We argue that each equilibrium corresponds to a distinct mode of interaction. For instance, in Fig. 1, yielding to the right or left are both acceptable equilibria, representing different Nash equilibria. Consequently, we propose that in order to capture different interaction modes, robots must reason about multiple interaction equilibria and select the one that aligns with the convention adopted by other agents. However, this is a non-trivial task as it requires the robot to find all local GNEs. To our best knowledge, there is no principled approach to compute multiple local Nash equilibria in constrained trajectory games other than exhaustive random initializations of the game solvers.
@@
In this paper, we tackle this challenge by introducing MultiNash-PF, a novel framework designed to efficiently compute multiple local GNEs in constrained trajectory games. This, in turn, enables robots to recognize and reason about multiple interaction modes. Our approach integrates 1) constrained potential games [13], a class of games for which local GNEs can be computed by solving a single constrained optimal control problem, and 2) constraint-aware implicit particle filtering which is a sample-efficient method recently applied to nonconvex trajectory optimization [14], to develop a principled approach for identifying various interaction outcomes.
@@
Using CPTGs, we can compute a local GNE by minimizing the potential function of the game, which is equivalent to solving a single constrained optimal control problem. We can solve the resulting optimal control problem using off-the-shelf optimization solvers. Then, we use a constraint-aware implicit particle filtering method to identify multiple local minima of the potential minimization problem. This method eliminates the need for random initialization of the optimization solver by efficiently obtaining coarse estimates of different local GNEs. Next, we use the identified minima as initialization for an optimization solver to refine our solution. The solver then computes all the different local GNEs of the original game, which correspond to multiple modes of interaction among the agents. Leveraging the sample efficiency of the implicit particle filtering method, MultiNash-PF can explore and identify multiple local GNEs faster than random initialization of optimization solvers. We further show the real-time application of MultiNash-PF in a human-robot navigation setting where the robot can reason about the existence of different local GNEs and plan its motion autonomously based on the modality the human prefers.
@@ II Related Works
Game Theoretic Planning. Game-theoretic planning has been widely used for modeling interactive domains. One of the foundational works in game theoretic motion planning was [15]. A wide range of methods have been proposed to compute different types of equilibria such as Stackelberg equilibria [16, 7], Nash equilibria, and generalized Nash equilibria through techniques such as best response dynamics [17, 18], terative linear-quadratic approximations [6, 10, 19], augmented Lagrangian-based solvers [20], and potential games-based methods  [13, 21, 11, 12, 22, 23]. However, the shortcomings of all these methods lie in the fact that they find only one Nash equilibrium of the underlying game and ignore the multi-modal nature of the problem that is crucial for coordination among agents. In [2], the authors recognize the existence of multiple equilibria and focus on inferring the equilibrium of other agents in the environment. However, their methodology relies on computing different approximate local Nash equilibria using rollouts of randomly selected initial strategies, which we show in this work performs poorly compared to our method.
Multiple Local Solutions to a Constrained Optimal Control Problem. Optimal control problems that are relevant to robotics are often nonconvex constrained optimization problems, which are solved using gradient-based methods such as interior-point algorithms [24] and sequential quadratic programming [25]. These methods, however, are limited to finding a single local solution per initialization. In the context of identifying all local GNEs, this limitation results in a significant computational burden as numerous initial guesses are required to explore the solution space. To address this, we depart from traditional approaches and adopt a Bayesian inference framework that enables simultaneous recovery of multiple local GNEs. Recent work [26, 27, 28] has demonstrated that constrained optimization problems can be reformulated as Bayesian inference tasks, enabling estimation algorithms to overcome the limitations of gradient-based solvers. In this setting, particle filtering methods [29, 30] provide a principled approach to capturing multimodal distributions. We leverage this by employing the implicit particle filter, which improves sample efficiency by focusing particles in high-probability regions of the posterior [31].
@@ III Constrained Potential Trajectory Games
Let $\mathcal{N}:=\{1,2,\ldots,N\}$ denote the set of agents’ indices. We assume that the game is played over a finite time horizon of $\tau\in\mathbb{N}$ time-steps, where $\mathbb{N}$ is the set of natural numbers. Let $x^{i}_{t}\in\mathbb{R}^{n_{i}}$ and $u_{t}^{i}\in\mathbb{R}^{m_{i}}$ denote the state and control input of the $i$-th agent at time $t$, respectively. We assume that the state of the $i$-th agent evolves as follows:
@@
We assume that each agent has to satisfy some constraints that couple different agents’ states and control inputs. Let $g:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{c}$ with $c\in\mathbb{N}$ denote a continuously differentiable function that defines the constraints on the joint state $x_{t}$ and joint control input $u_{t}$ at each time:
@@
where ${0}_{c}$ is a vector of zeros in $\mathbb{R}^{c}$. The aim of the constraint function $g$ is to capture individual agents’ strict preferences, such as each agent avoiding obstacles or avoiding collision with other agents. It should be noted that coupling amongst agents’ decisions is captured through constraints.
@@
where $\hat{x}_{t}^{i}\in\mathbb{R}^{n_{i}}$ is the reference state for agent $i$ at time $t$. Typically, the reference states for each agent is a path of minimum cost they would follow in the absence of other agents in the environment, e.g., a straight line connecting the start and goal location. Such a quadratic cost structure is a standard assumption in various motion planning domains [32]. Furthermore, we denote the joint reference state of all agents at time $t$ as $\hat{x}_{t}$. We denote by $\mathop{\rm blkdiag}(A_{1},\ldots,A_{n})$ the block-diagonal matrix formed by placing matrices $A_{1},\ldots,A_{n}$ on the diagonal with zeros elsewhere. Let $a_{0:\tau}:=\{a_{0},\ldots,a_{\tau}\}$ denote the collection of vectors $a_{t}$’s over the entire horizon $0\leq t\leq\tau$. We use $\{{x}_{0:\tau}^{i},{u}_{0:\tau}^{i}\}_{i=1}^{N}$ to denote the joint trajectory of all $N$ the agents at all times. Note that using the joint notation, we can equivalently write the joint trajectories of all agents as $\{{x}_{0:\tau},{u}_{0:\tau}\}$.
@@ III-A Local Generalized Nash Equilibrium Trajectories
Due to the interactive nature of the problem, generally, it is not possible for all agents to minimize their costs simultaneously while satisfying the constraints. Therefore, the interaction outcome is best captured by local GNE trajectories. With dynamics as (1) and constraints as (2), we denote the game by
@@
A set of joint trajectories $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}_{i=1}^{N}$ is a local GNE trajectory for $\mathcal{G}$ if for every agent $i\in\mathcal{N}$, the agent’s trajectory $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}$ is an optimal solution of the following optimization problem within a local neighborhood $\mathcal{D}\left(\{{x}^{\star}_{0:\tau},{u}^{\star}_{0:\tau}\},\epsilon\right)$ for some $\epsilon>0$:
@@
Intuitively, at a local GNE, no agent can reduce their cost function by independently changing their trajectory to any alternative feasible trajectory within the local neighborhood of the equilibrium trajectories. It is important to highlight that solving (1) requires solving a set of $N$ coupled constrained optimal control problems, which is often computationally expensive. Therefore, finding all the local GNEs of the game can be time-consuming when using traditional game-theoretic methods that aim to solve coupled optimal control problems.
However, an important property of our trajectory game is that it is a potential game, i.e., a GNE of the original game can be computed by optimizing a single constrained optimal control problem called Potential, as described in [11]. This reduces the computational cost of computing the GNEs of the game. In the following, we will prove that game $\mathcal{G}$ can be expressed as a constrained trajectory potential game and that any local minimizer of the potential function is a local GNE.
@@
A trajectory $\{{x}_{0:\tau}^{i^{\star}},{u}_{0:\tau}^{i^{\star}}\}_{i=1}^{N}$ is a local GNE if within a local neighborhood $\mathcal{D}\left(\{{x}^{\star}_{0:\tau},{u}^{\star}_{0:\tau}\},\epsilon\right)$, it is an optimal solution of
@@
Using the joint notations, we can re-write the sums in (6) as (1). Therefore, we have proven that $\mathcal{G}$ is a CPTG with potential function (1). Hence, using Theorem 2 from [11], any local solution of (1) will also be a local GNE of $\mathcal{G}$. ∎
Using Theorem 1, we want to find all local solutions of (1) that will also be local GNE trajectories as they correspond to different interaction modes and outcomes. We would like to acknowledge that finding local solutions of (1) will not necessarily give us all the GNEs of the original, as Theorem 1 provides only a sufficient condition for computing local GNEs. Furthermore, it is known that the problem of finding all Nash equilibria is a PPAD-complete problem even in the case of static games [33]. However, for practical purposes in multi-agent navigation, finding local solutions of (1) has proven to be enough for us.
The conventional approach to solving (1) relies on techniques such as interior point methods [24] or sequential quadratic programming [25]. Naively, the local solutions can be obtained by exhaustive initialization of random trajectories to an optimization solver, such as IPOPT [24], hoping to explore the solution space of local Nash equilibria. However, randomly exploring the nonconvex optimization landscape via gradient-based methods is computationally prohibitive and inefficient, as the same local equilibria can be arrived at from multiple random initializations. To address this, we utilize a Bayesian inference approach in the following section to efficiently recover all identified minima of (1).